Solve each of the following systems by substitution.
5x - 2y= -5
y - 5x = 3

8x - 4y = 16
y = 2x - 4


4x - 12y = 5
-x + 3y= -1

Answer by checkley75(3666)

Y-5X=3 OR Y=5X+3 NOW WE SUBSTITUTE (5X+3) FOR Y IN THE OTHER EQUATION & SOLVE FOR X:
5X-2(5X+3)=-5
5X-10X-6=-5
-5X=6-5
-5X=1
X=1/-5
X=-1/5 ANSWER.
Y-5*-1/5=3
Y+5/5=3
Y+1=3
Y=3-1
Y=2 ANSWER.
PROOF
5*-1/5-2*2=-5
-5/5-4=-5
-1-4=-5
-5=-5
I'LL LEAVE THE REST FOR YOU TO SOLVE USING THE SAME PROCEDURE!